OpenStudy (anonymous):

Use the distributive property to simplify the expression below: -5x-10/-5

OpenStudy (anonymous):

Take -5 common from numerator, -5(x + 2)/(-5) = x + 2 is the answer..

OpenStudy (anonymous):

-5x-10 -5(x + 2) ------- = ------------ = x + 2 -5 -5

OpenStudy (maheshmeghwal9):